OpenStudy (anonymous):

divide. (x^2+2x+1)/(x-2)/(x^2-1)/(x^2-4)

OpenStudy (anonymous):

X=-1

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

if x=-1, you have a point of discontinuity because then part of the problem is dividing by zero which is a no-no. I also don't think that's the goal of this problem.

OpenStudy (anonymous):

well i did not lie X=-1

OpenStudy (callisto):

First, factorize the expressions. For (x^2+2x+1), use the identity \((a+b)^2= a^2 + 2ab+ b^2\) to factorize it. In this case, a is x, b is 1. So, x^2 + 2x +1 = (x+1)^2 For (x^2-1), use the identity \(a^2-b^2 = (a+b)(a-b)\). In this case, a is x, b is 1. x^2 - 1 = (x-1)(x+1) Similar for (x^2-4) x^2 - 4 = x^2 - 2^2 = (x-2)(x+2) Now, for your question, \[\frac{\frac{(x^2+2x+1)}{(x-2)}}{\frac{(x^2-1)}{(x^2-4)}}\]\[=\frac{\frac{(x+1)^2}{(x-2)}}{\frac{(x+1)(x-1))}{(x-2)(x+2)}}\]\[=\frac{(x+1)^2}{(x-2)}\times \frac{(x-2)(x+2)}{(x+1)(x-1)}\] Cancel the common factor to get the answer..

OpenStudy (anonymous):

(x+1)(x-2)/x-1

OpenStudy (anonymous):

x^2+2+1/(x-2)(x2-4)(x^2-1)

OpenStudy (callisto):

Hmm....Perhaps not yet finished... (x+1)(x-2) = x^2 -2x + x -2 = x^2 - x -2 Now, use synthetic division / long division.. x ---------------- x-1 ) x^2 - x - 2 x^2 - x --------------- -2 You can get the quotient and remainder..

OpenStudy (anonymous):

thank you everyone good night

OpenStudy (anonymous):

your welcome c(;

OpenStudy (anonymous):

I just aced my test the correct answer is (x+1)(x+2)/x-1